OpenStudy (anonymous):

find inverse y=(3^x)/(2+3^x)

OpenStudy (loser66):

wow.... but solvable step 1: switch the role of x and y \[x = \dfrac{3^y}{2+3^y}\] step2, solve for y

OpenStudy (loser66):

done

OpenStudy (anonymous):

well yes u need to seperate y by itself though

OpenStudy (anonymous):

i got that. how do u solve for "Y"???

OpenStudy (loser66):

mmmmhhhh.... I know just that, the leftover is your duty. and the sentence "I got that" sounds familiar to me, hehehe

OpenStudy (rational):

you familiar wid logs and stuff ?

OpenStudy (anonymous):

no im serious. i got that but i have questions for just the next step. & yes i am familiar

OpenStudy (rational):

good, start by crossmultiplying

OpenStudy (rational):

\[x = \dfrac{3^y}{2+3^y}\] \[x(2+3^y) = 3^y\] \[2x+x3^y = 3^y\] \[2x= 3^y(1-x)\] \[\dfrac{2x}{1-x}= 3^y\]

OpenStudy (rational):

use ur log properties to separate y

OpenStudy (anonymous):

do you know what log properties to use?

OpenStudy (anonymous):

ok wait i think i got it. is it this?

OpenStudy (loser66):

nope

OpenStudy (anonymous):

why?

OpenStudy (loser66):

continue the stuff of @rational apply the equivalent of exponential and log

OpenStudy (anonymous):

no... work through it again

OpenStudy (rational):

thats really a crazy abusive way of using(inventing) log properties lol :P

OpenStudy (loser66):